Accelerator Update for June 1, 2009 - June 3, 2009 |
Monday June 1 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (144.4E10), with the Antiproton Source stacking (145.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson, MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. Operations established store 7101 at 8:15 AM with an average initial luminosity of 260E30. Linac personnel addressed problems with a low-energy RF station (LRF4) and a Klystron RF station (KRF3), which caused a small amount of downtime. MI studies began following the Linac downtime. The Pelletron experienced high voltage (HV) regulation problems and required time to condition the accelerating structure. This lasted until about 9:30 PM and delayed the store termination. As a result antiproton transfers to the Recycler were halted until after HEP shot setup. Operations terminated store 7101 at 10:50 PM and put the TeV into HEP shot setup. Pbar transferred antiprotons to the Recycler when appropriate. Tuesday June 2 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (187.1E10), with Pbar stacking (120.8E10), and with NuMI and MiniBooNE taking beam. Operations established store 7102 at 12:45 AM with an average initial luminosity of 270E30. Pelletron HV regulation problems resurfaced around 4 AM. Operations contacted Pelletron personnel to assist in its recovery; the accelerating structure was conditioned for a short time while waiting for their arrival. Efforts to stabilize the HV continued through out the rest of the midnight shift. Antiproton transfers to the Recycler were halted and the stash remained at ~250E10. A Pelletron access was required to resolve the HV regulation problems. Efforts to prepare the SF6 gas transfer out of the Pelletron continued through out the day shift. The Pbar cryo system dry engine flywheel failed at 2:15 PM requiring repairs and a halt to stacking. Booster, MI, Pbar, and TeV personnel conducted various machine studies through out most of the day shift. Operations disabled the MTest beam at 2:30 PM; this ended T979’s current run. A different MTest experiment (T988) began preparing move in. The Pbar dry engine repairs were completed and beam was reestablished at around 7 PM. The SF6 gas transfers were completed and the Pelletron was ready for access at 7 PM. There were no antiproton transfers to the Recycler due to the Pelletron repairs. Wednesday June 3 The midnight shift began with Operations monitoring store 7102 (46.47E30) and the stash (241.9E10), with Pbar stacking (196.3E10), and with NuMI and MiniBooNE taking beam. Pelletron repairs continued through out most of the midnight shift. Final tests were completed and Pelletron personnel begin transferring gas back into the Pelletron by 4 AM. Ops monitored the 240E10 stash and the rather sizable stack. No antiproton transfers were made to the Recycler due to the Pelletron repairs. The Plans for Wednesday The plans for today are to finish the Pelletron repairs and resume normal stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 06/03/2009 email Fermilab |
| Security, Privacy, Legal | |